OpenStudy (anonymous):

Will Medal or Fan! What is the value of b? 11 + 4b =8b - 13 B = _____

OpenStudy (anonymous):

b= 6

OpenStudy (midhun.madhu1987):

8b - 4b = 11 + 13 4b = 24 so b= ?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@KissMyAxe

OpenStudy (anonymous):

@Paradise9541 its alright :)